The International Association of Bulk Marine Carriers (tramp steamers) is unhappy with the current pattern of the boom and bust price cycles in their industry. They want to form a “shipping conference” like the liner carriers (container ships) to set prices and output. They have asked you to examine the feasibility of forming a cartel, given that international law does not prevent them from creating such organizations.

a. Draw an appropriate economic model(s) and explain whether or not, this industry has the conditions for a cartel organization to be a successful.

b. The bulk marine carriers’ customers are raw material shippers (minerals, coal, grain, etc.) that experience demand shifts as the world economy expands and contracts. Would this make a cartel more difficult to manage? Explain.
Answer #1

a. Cartels as we know are collusive agreements created for avoiding uncertainty arising from oligopolistic interdependence. Cartels are of two types.

A. Cartels aiming at joint- profit maximisation.

B. Market- sharing cartels.

The former is that there is a central agency to whom all the members of the cartel gives the authority to decide the total quantity, price to gain maximum profit, allocation of production among the members of the cartel, and also the distribution of the maximum joint profit among the participating members. Also the central agency has access to the cost figures of the members. The market MC curve is derived from the horizontal summation of the member MC curves and the central agency sets the price where this MC equals the MR of the industry.

Presenting the model:

the market demand: P = 100 - 0.5(X) = 100 - 0.5(X1+X2)

If we take there are only two firms in International Association of Bulk Marine Carriers, then the costs of both ofthem are:

C1 = 5X1 and C2 = 0.5X2

The central agency aims at :

\prod= \prod1+\prod2

where \prod1 = Revenue 1- Cost 1 and \prod2 = Revenue 2- Cost 2

\prod= [100-0.5(X1+X2)](X1+X2)-5X1-0.5X2

= 95X1-100X2-0.5X12 -X22-X1X2

Setting the partial derivatives:

ОПох1 = 85 -X1 -X2 = 0

ОПох? = 100- X1 -2X2 = 0

Solving for X1 and X2

X1=90 X2=5

Substituting,

P = 100-0.5 (X1+X2) = 52.5

Joint profit:

\prod = 95 (90) - 100(5) - 0.5 (90)2 -(5)2 -(90)(5) = 4525

The latter, market sharing cartels member firms agree on a common price at which each of the can sell any quantity demanded. And the price can be set by bargaining, with the low-cost firms for a low price and the high cost firms with high price, but they are free to vary the style of the product or their selling activities. Under the market sharing there is another type of cartels wherein the sharing of market is based on quotas, that is agreement on the quantity each member can sell at an agreed price.

The International Association of Bulk Marine Carriers can form a good cartel and take the market sharing cartel form so that they can secure themselves from the frequent change in prices and quantities.

b. The demand shifts cannot make a cartel difficult because they are collusive agreements. So what ever be the changes in the world market, this is effected for the members eqully. So it is not much difficult to deal with.